# On the gcd graphs over polynomial rings and related topics

### v1

@@ 7. Perfect gcd-graphs
In [1, Theorem 9.5], the authors classify all perfect unitary Cayley graphs associated with finite commutative rings. Specifically, for a ring $R=R_{1}\times R_{2}\times\cdots\times R_{t}$ where $R_{i}$ are finite local rings such that $|R_{1}|\leq|R_{2}|\leq\ldots\leq|R_{t}|$, the unitary Cayley graph on $R$ is perfect if and only if one of the following conditions hold
@@
    The residue field of $R_{1}$ is $2$. In this case, $G_{R}$ is bipartite and hence perfect.
@@
A direct consequence of [1, Theorem 9.5] for the unitary Cayley graph over $\operatorname{\mathbb{Z}}$ is the following.
@@
The unitary Cayley graph on $\operatorname{\mathbb{Z}}/N$ is perfect if and only if one of the following conditions holds.
@@
The unitary Cayley graph $G_{f}(\{1\})$ on $\operatorname{\mathbb{F}}_{q}[x]/f$ is perfect if and only if one of the following conditions holds.
@@
We will use 7.2 together with patterns of $G_{f}(D)$ discovered through our experiments in Sagemath and the Python package Networkx in order to find some sufficient conditions for $G_{f}(D)$ to be non-perfect. In particular, we will exploit the fact that certain induced subgraphs of $G_{f}(D)$ are naturally isomorphic to the unitary Cayley graphs on some quotient rings of $R=\operatorname{\mathbb{F}}_{q}[x]/f.$ More precisely, by 5.4, we have the following observation.
@@
$\operatorname{\mathbb{F}}_{q}[x]/f\cong\operatorname{\mathbb{F}}_{q}[x]/f_{1}\times\operatorname{\mathbb{F}}_{q}[x]/f_{2}\times\operatorname{\mathbb{F}}_{q}[x]/f_{3}.$
Under this isomorphism, we can identify $V(G_{f}(D))$ as the set of all triples $(a_{1},a_{2},a_{3})$ where $a_{i}\in\operatorname{\mathbb{F}}_{q}[x]/f_{i}$ for $1\leq i\leq 3$ Furthermore, two vertices $(a_{1},a_{2},a_{3})$ and $(b_{1},b_{2},b_{3})$ are adjacent if and only if there exists an index $i\in\{1,2,3\}$ such that $a_{i}=b_{i}$ and $(a_{j}-b_{i})\in(\operatorname{\mathbb{F}}_{q}[x]/f_{j})^{\times}$ if $j\neq i$. Using Sagemath, we can find the following induced $5$-cycle in $G_{f}(D)$
@@
$\operatorname{\mathbb{F}}_{q}[x]/g\cong\operatorname{\mathbb{F}}_{q}[x]/h_{1}\times\operatorname{\mathbb{F}}_{q}[x]/h_{2}\times\operatorname{\mathbb{F}}_{q}[x]/f_{2}.$
Under this isomorphism, we can identify $\operatorname{\mathbb{F}}_{q}[x]/g$ with the set of tuples $(a_{1},a_{2},a_{3})$ such that $a_{1}\in\operatorname{\mathbb{F}}_{q}[x]/h_{1},a_{2}\in\operatorname{\mathbb{F}}_{q}[x]/h_{2},a_{3}\in\operatorname{\mathbb{F}}_{q}[x]/f_{2}.$ Furthermore, two vertices $(a_{1},a_{2},a_{3})$ and $(b_{1},b_{2},b_{3})$ are adjacent if and only if one of the following conditions happens
@@
    $a_{1}=b_{1},a_{2}=b_{2}$ and $a_{3}-b_{3}\in(\operatorname{\mathbb{F}}_{q}[x]/f_{2})^{\times}$,
@@
    $a_{3}=b_{3}$ and $a_{i}-b_{i}\in(\operatorname{\mathbb{F}}_{q}[x]/h_{i})^{\times}$ for $i\in\{1,2\}.$
@@
Curious readers might wonder why we choose a $7$-cycle in the proof of 7.5 instead of choosing a $5$-cycle as in the proof of 7.4. The reason is that when $\omega(f_{1})=2,\omega(f_{2})=1$ and $f=f_{1}f_{2}$, our code cannot find $5$-cycle in $G_{f}(D).$ It seems interesting to investigate whether this is always the case.
@@
  1. (1)
    If $f=f_{1}f_{2}$ where both $f_{1},f_{2}$ are irreducible then $G_{f}(D)$ is isomorphic to the complement of $G_{f}(\{1\})$ which is perfect by 7.2. Therefore, $G_{f}(D)$ is perfect as well.
  2. (2)
    On the other hand, if either $f_{1}$ or $f_{2}$ is reducible, our code can always find a $7$-cycle in $G_{f}(D).$ Unfortunately, we cannot find a universal pattern in this case. It would be quite interesting to solve this puzzle completely.
@@ 8. Induced subgraphs of gcd-graphs
A theorem of Erdős and Evans (see [11]) says that every graph $G$ is an induced subgraph of the unitary Cayley graph on $\operatorname{\mathbb{Z}}/n$ for some squarefree $n.$ Using this result, it is shown in [16] that for a given finite graph $G$ and a finite field $F$, $G$ is an induced subgraph of the unitary graph of a matrix algebra $M_{d}(F)$ for some value of $d$ (they also provide some precise upper-bound on $d$ when $G$ is the complete graph $K_{m}$). In light of these results, it seems interesting to ask whether a graph $G$ can be realized as an induced subgraph of $G_{f}(D)$ for some choice of $f,D$ and $\operatorname{\mathbb{F}}_{q}.$ Under some rather mild conditions, the answer is yes as we will show below. First, we introduce the following observation.
@@
If $G$ is an induced subgraph of $H$, then $G$ is also an induced subgraph of the product $H\times K_{|G|}$ where $K_{|G|}$ is the complete graph on $|G|$-nodes.
@@
Let $G$ be a fixed graph and $q$ a fixed prime power. There exists a positive integer $r$ such that for each $d\geq r$, we can find $F_{d}\in\operatorname{\mathbb{F}}_{q}[x]$ satisfying the following conditions
@@
By Erdős and Evans’s theorem [11], there exists a squarefree integer $n\in\operatorname{\mathbb{N}}$ such that $G$ is an induced subgraph of the unitary graph $G_{n}$ on $\operatorname{\mathbb{Z}}/n.$ Let $r=\omega(n)$; i.e, $n=p_{1}p_{2}\ldots p_{r}$ be the prime factorization of $n$. Then
@@
For each $i\in\operatorname{\mathbb{N}}$ and $m_{i}\in\operatorname{\mathbb{N}}$, there exists a polynomial $h_{i}$ of degree $m_{i}$ such that $h_{i}$ is irreducible over $\operatorname{\mathbb{F}}_{q}[x].$ We then see that $G_{h_{i}}(\{1\})$ is isomorphic to the complete graph $K_{q^{m_{i}}}.$ Let us choose $m_{i}$ so that $q^{m_{i}}>n$ for all $i$. For each $d\geq r$, let $F_{d}=h_{1}h_{2}\ldots h_{d}.$ We then have
@@
Since $q^{m_{i}}>n$, we know that $G_{n}$ is an induced subgraph of $G_{F_{r}}(\{1\}).$ By 8.1 , $G$ is an induced subgraph of $G_{F_{d}}(\{1\})$ as well.
@@
Let $I$ be the ideal generated by $h_{0}$ in $\operatorname{\mathbb{F}}_{q}[x]/f.$ By 5.4 the induced graph on $I$ is naturally isomorphic to the unitary Cayley graph $G_{h}(\{1\})$. This shows that $G$ is an induced subgraph of $G_{f}(D).$
@@
    $G$ is an induced subgraph of $G_{f}(D)$ where $D=\{f_{1},f_{2},\ldots,f_{k}\}$.
In general, the answer is no. There seem to be some subtle constraints. We discuss here a particle one. Let $g=\textrm{lcm}(f_{1},f_{2},\ldots,f_{k})$ and assume further that $g\neq f_{i}$ for all $1\leq i\leq k$. If such $f$ exists, then $g\mid f$ and there is a canonical map
$\Phi:\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{F}}_{q}[x]/g.$
If we take a subset $S\subset\operatorname{\mathbb{F}}_{q}[x]/f$ such that $|S|>|\operatorname{\mathbb{F}}_{q}[x]/g|$, then there exists $a,b\in S$ such that $a\neq b$ and $\Phi(a)=\Phi(b).$ Consequently, $\Phi(a-b)=0$ or $g\mid a-b.$ By definition, $(a,b)\not\in E(G_{f}(D)).$ Consequently, we have the following upper bound for the clique number of $G_{f}(D)$
$\omega(G_{f}(D))\leq|\operatorname{\mathbb{F}}_{q}[x]/g|.$
This shows that if $\omega(G)>|\operatorname{\mathbb{F}}_{q}[x]/g|$ then $G$ cannot be an induced subgraph of $G_{f}(D).$
@@
Let $m$ be a positive integer such that $q^{m}>n$ and $g=\textrm{lcm}(f_{1},f_{2},\ldots,f_{k})\in\operatorname{\mathbb{F}}_{q}[x].$ By Galois theory, there exists a polynomial $h\in\operatorname{\mathbb{F}}_{q}[x]$ with at least $r$ distinct irreducible factors and $\gcd(h,g)=1.$ Let $f=hg$. Let $I$ be the ideal in $\operatorname{\mathbb{F}}_{q^{m}}[x]/f$ generated by $f_{1}.$ Then, by 5.4, the induced subgraph on $I$ is isomorphic to the unitary Cayley graph $G_{\bar{f}}(\{1\})$ where $\bar{f}=f/f_{1}.$ Suppose that $\bar{f}=g_{1}^{a_{1}}g_{2}^{a_{2}}\ldots g_{t}^{a_{t}}$ be the factorization of $f/f_{1}$ over $\operatorname{\mathbb{F}}_{q^{m}}[x].$ Then, by the choice of $f$, $t\geq r.$ By the Chinese remainder theorem, we know that $\operatorname{\mathbb{F}}_{q^{m}}^{t}$ is a subring of $\operatorname{\mathbb{F}}_{q^{m}}[x]/\bar{f}.$ Therefore, $G_{\operatorname{\mathbb{F}}_{q^{m}}^{t}}=\prod_{i=1}^{t}K_{q^{m}}$ is an induced subgraph of $G_{\bar{f}}(\{1\})$. Since $t\geq r$ and $q^{m}\geq n$, by 8.1 $G_{n}$ is an induced subgraph of $G_{\operatorname{\mathbb{F}}_{q^{m}}^{t}}$. Consequently, $G_{n}$ is an induced subgraph of $G_{f}(D)$ as well. ∎
@@ Acknowledgements
We thank Professor Torsten Sander for the enlightening correspondence about gcd-graphs and related arithmetic graphs. We are grateful to Professor Emanuel Kolwaski for some helpful comments on Fekete polynomials and for writing the book [19]. We thank Professor Ki-Bong Nam for his interest in this topic and his feedback on the first draft. Finally, we thank Minh-Tam Trinh for his discussions, encouragement, and support at the initial stage of this work.

### v2

@@ 7. Perfect gcd-graphs
In [1, Theorem 9.5], the authors classify all perfect unitary Cayley graphs associated with finite commutative rings. Specifically, for a ring $R=R_{1}\times R_{2}\times\cdots\times R_{t}$ where $R_{i}$ are finite local rings such that $|R_{1}|\leq|R_{2}|\leq\cdots\leq|R_{t}|$, the unitary Cayley graph on $R$ is perfect if and only if one of the following conditions hold:
@@
    The residue field of $R_{1}$ is $2$. In this case, $G_{R}$ is bipartite and hence perfect;
@@
[…]
@@
The unitary Cayley graph on $\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}$ is perfect if and only if one of the following conditions holds.
@@
The unitary Cayley graph $G_{f}(\{1\})$ on $\operatorname{\mathbb{F}}_{q}[x]/(f)$ is perfect if and only if one of the following conditions holds.
@@
We will use 7.2 together with patterns of $G_{f}(D)$ discovered through our experiments in Sagemath and the Python package Networkx in order to find some sufficient conditions for $G_{f}(D)$ to be non-perfect. In particular, we will exploit the fact that certain induced subgraphs of $G_{f}(D)$ are naturally isomorphic to the unitary Cayley graphs on some quotient rings of $R=\operatorname{\mathbb{F}}_{q}[x]/(f).$ More precisely, by 5.4, we have the following observation.
@@
$\operatorname{\mathbb{F}}_{q}[x]/(f)\cong\operatorname{\mathbb{F}}_{q}[x]/(f_{1})\times\operatorname{\mathbb{F}}_{q}[x]/(f_{2})\times\operatorname{\mathbb{F}}_{q}[x]/(f_{3}).$
Under this isomorphism, we can identify $V(G_{f}(D))$ as the set of all triples $(a_{1},a_{2},a_{3})$ where $a_{i}\in\operatorname{\mathbb{F}}_{q}[x]/(f_{i})$ for $1\leq i\leq 3.$ Furthermore, two vertices $(a_{1},a_{2},a_{3})$ and $(b_{1},b_{2},b_{3})$ are adjacent if and only if there exists an index $i\in\{1,2,3\}$ such that $a_{i}=b_{i}$ and $(a_{j}-b_{i})\in(\operatorname{\mathbb{F}}_{q}[x]/(f_{j}))^{\times}$} if $j\neq i$. Using Sagemath, we can find the following induced $5$-cycle in $G_{f}(D)$
@@
$\operatorname{\mathbb{F}}_{q}[x]/(g)\cong\operatorname{\mathbb{F}}_{q}[x]/(h_{1})\times\operatorname{\mathbb{F}}_{q}[x]/(h_{2})\times\operatorname{\mathbb{F}}_{q}[x]/(f_{2}).$
Under this isomorphism, we can identify $\operatorname{\mathbb{F}}_{q}[x]/(g)$ with the set of tuples $(a_{1},a_{2},a_{3})$ such that $a_{1}\in\operatorname{\mathbb{F}}_{q}[x]/(h_{1}),a_{2}\in\operatorname{\mathbb{F}}_{q}[x]/(h_{2}),a_{3}\in\operatorname{\mathbb{F}}_{q}[x]/(f_{2}).$ Furthermore, two vertices $(a_{1},a_{2},a_{3})$ and $(b_{1},b_{2},b_{3})$ are adjacent if and only if one of the following conditions happens
@@
    $a_{1}=b_{1},a_{2}=b_{2}$ and $a_{3}-b_{3}\in(\operatorname{\mathbb{F}}_{q}[x]/(f_{2}))^{\times}$,
@@
    $a_{3}=b_{3}$ and $a_{i}-b_{i}\in(\operatorname{\mathbb{F}}_{q}[x]/(h_{i}))^{\times}$ for $i\in\{1,2\}.$
@@
Curious readers might wonder why we choose a $7$-cycle in the proof of 7.5 instead of choosing a $5$-cycle as in the proof of 7.4. The reason is that when $\omega(f_{1})=2,\omega(f_{2})=1$ and $f=f_{1}f_{2}$, our code cannot find a $5$-cycle in $G_{f}(D).$ It seems interesting to investigate whether this is always the case.
@@
(1) If $f=f_{1}f_{2}$ where both $f_{1},f_{2}$ are irreducible then $G_{f}(D)$ is isomorphic to the complement of $G_{f}(\{1\})$ which is perfect by 7.2. Therefore, $G_{f}(D)$ is perfect as well.
(2) On the other hand, if either $f_{1}$ or $f_{2}$ is reducible, our code can always find a $7$-cycle in $G_{f}(D).$ Unfortunately, we cannot find a universal pattern in this case. It would be quite interesting to solve this puzzle completely.
@@ 8. Induced subgraphs of gcd-graphs
A theorem of Erdős and Evans (see [12]) says that every graph $G$ is an induced subgraph of the unitary Cayley graph on $\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}$ for some squarefree $n.$ Using this result, it is shown in [17] that for a given finite graph $G$ and a finite field $F$, $G$ is an induced subgraph of the unitary graph of a matrix algebra $M_{d}(F)$ for some value of $d$ (they also provide some precise upper-bound on $d$ when $G$ is the complete graph $K_{m}$). In light of these results, it seems interesting to ask whether a graph $G$ can be realized as an induced subgraph of $G_{f}(D)$ for some choice of $f,D$ and $\operatorname{\mathbb{F}}_{q}.$ Under some rather mild conditions, the answer is yes as we will show below. For the related definitions we refer the reader to Section 2. First, we introduce the following observation.
@@
If $G$ is an induced subgraph of $H$, then $G$ is also an induced subgraph of the tensor product $H\times K_{|G|}$ where $K_{|G|}$ is the complete graph on $|G|$-nodes.
@@
Let $G$ be a fixed graph and $q$ a fixed prime power. There exists a positive integer $r$ such that for each $d\geq r$, there exists $F_{d}\in\operatorname{\mathbb{F}}_{q}[x]$ satisfying the following conditions
@@
By Erdős and Evans’s theorem [12], there exists a squarefree integer $n\in\operatorname{\mathbb{N}}$ such that $G$ is an induced subgraph of the unitary graph $G_{n}$ on $\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}.$ Let $r=\omega(n)$; i.e, $n=p_{1}p_{2}\ldots p_{r}$ be the prime factorization of $n$. Then
@@
For each $i\in\operatorname{\mathbb{N}}$ and $m_{i}\in\operatorname{\mathbb{N}}$, by Galois theory for finite fields, there exists a polynomial $h_{i}$ of degree $m_{i}$ such that $h_{i}$ is irreducible over $\operatorname{\mathbb{F}}_{q}[x].$ We then see that $G_{h_{i}}(\{1\})$ is isomorphic to the complete graph $K_{q^{m_{i}}}.$ Let us choose $m_{i}$ so that $q^{m_{i}}>n$ for all $i$. For each $d\geq r$, let $F_{d}=h_{1}h_{2}\ldots h_{d}.$ We then have
@@
Since $q^{m_{i}}>n$, we know that $G_{n}$ is an induced subgraph of $G_{F_{r}}(\{1\}).$ By 8.1 , $G$ is an induced subgraph of $G_{F_{d}}(\{1\})$ as well. ∎
We obtain the following result which says in particular that every graph is an induced subgraph of a gcd-graph.
@@
Let $I$ be the ideal generated by $h_{0}$ in $\operatorname{\mathbb{F}}_{q}[x]/(f).$ By 5.4 the induced graph on $I$ is naturally isomorphic to the unitary Cayley graph $G_{h}(\{1\})$. This shows that $G$ is an induced subgraph of $G_{f}(D).$
@@
    $G$ is an induced subgraph of $G_{f}(D)$ where $D=\{f_{1},f_{2},\ldots,f_{k}\}$?
In general, the answer is no. There seem to be some subtle constraints. We discuss here a particular one. Let $g=\textrm{lcm}(f_{1},f_{2},\ldots,f_{k})$ and assume further that $g\neq f_{i}$ for all $1\leq i\leq k$. If such $f$ exists, then $g\mid f$ and there is a canonical map
$\Phi\colon\operatorname{\mathbb{F}}_{q}[x]/(f)\to\operatorname{\mathbb{F}}_{q}[x]/(g).$
If we take a subset $S\subset\operatorname{\mathbb{F}}_{q}[x]/(f)$ such that $|S|>|\operatorname{\mathbb{F}}_{q}[x]/(g)|$, then there exists $a,b\in S$ such that $a\neq b$ and $\Phi(a)=\Phi(b).$ Consequently, $\Phi(a-b)=0$ or $g\mid a-b.$ By definition, $(a,b)\not\in E(G_{f}(D)).$ Consequently, we have the following upper bound for the clique number of $G_{f}(D)$
$\omega(G_{f}(D))\leq|\operatorname{\mathbb{F}}_{q}[x]/(g)|.$
This shows that if $\omega(G)>|\operatorname{\mathbb{F}}_{q}[x]/(g)|$ then $G$ cannot be an induced subgraph of $G_{f}(D).$
@@
Let $m$ be a positive integer such that $q^{m}>n$ and $g=\textrm{lcm}(f_{1},f_{2},\ldots,f_{k})\in\operatorname{\mathbb{F}}_{q}[x].$ By Galois theory, there exists a polynomial $h\in\operatorname{\mathbb{F}}_{q}[x]$ with at least $r$ distinct irreducible factors and $\gcd(h,g)=1.$ Let $f=hg$. Let $I$ be the ideal in $\operatorname{\mathbb{F}}_{q^{m}}[x]/(f)$ generated by $f_{1}.$ Then, by 5.4, the induced subgraph on $I$ is isomorphic to the unitary Cayley graph $G_{\bar{f}}(\{1\})$ where $\bar{f}=f/f_{1}.$ Suppose that $\bar{f}=g_{1}^{a_{1}}g_{2}^{a_{2}}\ldots g_{t}^{a_{t}}$ be the factorization of $f/f_{1}$ over $\operatorname{\mathbb{F}}_{q^{m}}[x].$ Then, by the choice of $f$, $t\geq r.$ By the Chinese remainder theorem, we know that $\operatorname{\mathbb{F}}_{q^{m}}^{t}$ is a subring of $\operatorname{\mathbb{F}}_{q^{m}}[x]/\bar{f}.$ Therefore, $G_{\operatorname{\mathbb{F}}_{q^{m}}^{t}}=\prod_{i=1}^{t}K_{q^{m}}$ is an induced subgraph of $G_{\bar{f}}(\{1\})$. Since $t\geq r$ and $q^{m}\geq n$, by 8.1 $G_{n}$ is an induced subgraph of $G_{\operatorname{\mathbb{F}}_{q^{m}}^{t}}$. Consequently, $G_{n}$ is an induced subgraph of $G_{f}(D)$ as well. ∎
@@ Acknowledgements
We thank Professor Torsten Sander for the enlightening correspondence about gcd-graphs and related arithmetic graphs. We are grateful to Professor Emmanuel Kolwaski for some helpful comments on Fekete polynomials and for writing the book [20]. We thank Professor Ki-Bong Nam for his interest in this topic and his feedback on the first draft. Finally, we thank Minh-Tam Trinh for his discussions, encouragement, and support at the initial stage of this work. We are also grateful to the referee for their comments and valuable suggestions which we have used to improve our exposition.